# Gaussian Copula Models for Nonignorable Missing Data Using Auxiliary Marginal Quantiles

### v1

@@ Abstract
We present an approach for modeling and imputation of nonignorable missing data under Gaussian copulas. The analyst posits a set of quantiles of the marginal distributions of the study variables, for example, reflecting information from external data sources or elicited expert opinion. When these quantiles are accurately specified, we prove it is possible to consistently estimate the copula correlation and perform multiple imputation in the presence of nonignorable missing data. We develop algorithms for estimation and imputation that are computationally efficient, which we evaluate in simulation studies of multiple imputation inferences. We apply the model to analyze associations between lead exposure levels and end-of-grade test scores for 170,000 students in North Carolina. These measurements are not missing at random, as children deemed at-risk for high lead exposure are more likely to be measured. We construct plausible marginal quantiles for lead exposure using national statistics provided by the Centers for Disease Control and Prevention. Complete cases and missing at random analyses appear to underestimate the relationships between certain variables and end-of-grade test scores, while multiple imputation inferences under our model support stronger, adverse associations between lead exposure and educational outcomes.
Keywords: Bayesian; imputation; MNAR; data integration; nonresponse
@@ 1 Introduction
The Gaussian copula is a flexible joint distribution for multivariate data. The model can characterize complex dependencies while also capturing non-Gaussian marginal distributions. Methodological advances have made the Gaussian copula compatible with mixed data types (Hoff, , 2007; Feldman and Kowal, , 2022), increased its scalability to high-dimensional variable sets (Murray et al., , 2013), and improved its ability to capture non-linearity and interactions (Feldman and Kowal, , 2024). Because of these appealing features, it has been deployed in numerous applications including, for example, in economics and finance (Fan and Patton, , 2014), marketing and management (Eckert and Hohberger, , 2023), and political science (Chiba et al., , 2015).
@@
The Gaussian copula can be readily implemented when data have missing values. Indeed, researchers (e.g., Käärik and Käärik, , 2010; Di Lascio et al., , 2015; Zhao and Udell, , 2020; Hollenbach et al., , 2021; Hoff, , 2022; Christoffersen et al., , 2023) have suggested Gaussian copulas as multivariate, joint modeling engines for (multiple) imputation. Existing methods, however, tend to assume that missingness among the study variables is missing completely at random (MCAR) or missing at random (MAR) (Rubin, , 1976). We are not aware of methodology for itemwise nonignorable missingness with Gaussian copula models.
@@
In this article, we develop a framework to handle nonignorable missing data in Gaussian copula models. Our general strategy is to specify a Bayesian Gaussian copula model that includes both the study variables and the missingness indicators, coupled with auxiliary information on the marginal distributions of the study variables. Using this strategy, we prove that analysts need to know only a small, arbitrary set of auxiliary marginal quantiles to enable consistent estimation of the copula correlation; entire distributions are not necessary. These results hold for a version of an additive nonignorable (AN) missingness mechanism (Hirano et al., , 2001; Sadinle and Reiter, , 2019). This mechanism allows the reason for missingness in a variable potentially to depend on the value of the variable itself. We develop algorithms for estimating the copula correlation that result in significant computational gains relative to alternative copula models, which allows fitting to scale to data with (moderately) large sample sizes in reasonable time using typical computational setups. We also present strategies for estimating other quantiles of the marginal distributions beyond those in the auxiliary information. Our approach provides uncertainty quantification for the unknown marginals, which we utilize for multiple imputation (Rubin, , 1987).
@@
We develop the methodology for settings where marginal quantiles for variables subject to missingness are available in external data sources. For example, previewing the application in Section 5, suppose an analysis involving health measurements suffers from nonignorable missingness, e.g., people likely to have unconcerning values are not measured. For many health measurements, there exists information on marginal quantiles from national benchmark surveys or administrative databases. We seek to use the auxiliary quantiles to adjust for nonignorable missingness in the analysis at hand. When auxiliary quantiles are not available or known precisely, analysts can posit different plausible marginal quantiles and assess the sensitivity of ultimate analyses under those specifications, e.g., via multiple imputation analyses. We note that the use of known marginal distributions for imputation of nonignorable missing data also has been proposed for categorical data models (Pham et al., , 2018; Akande et al., , 2021; Deng et al., , 2013; Si et al., , 2015, 2016; Tang et al., , 2024), but not for continuous and mixed variables like we do here.
@@
We apply the methodology in data comprising health, socioeconomic, demographic and educational measurements collected on over 170,000 North Carolina children. Among the study variables are end-of-grade math and reading test scores and a measure of lead exposure, the latter of which is subject to abundant missingness that is likely nonignorable. The state requires children at high risk of lead exposure to be measured, but not children at low risk of lead exposure. To inform the imputation of the missing values, we leverage marginal quantiles on lead exposure published by the Centers for Disease Control and Prevention. Using this information, we find lead exposure apparently is more adversely associated with test scores than suggested from complete cases or MAR analyses.
@@
The remainder of this article is organized as follows. In Section 2, we present the Gaussian copula model with nonignorable missing data and information on marginal distributions. Here we present two theoretical results, namely (i) that fully specified marginal distributions provide posterior consistency of the copula correlation in the presence of nonignorable missingness, and (ii) a derivation that the Gaussian copula implies a version of the AN missingness mechanism. In Section 3, we modify the copula model and develop imputation strategies for settings where the auxiliary marginal information constitutes a specified set of marginal quantiles. Here we present our main theorem: under the AN missingness mechanism, this limited auxiliary information still allows consistent estimation of the copula correlation. In Section 4, we present simulations studying the effect of differing amounts of auxiliary quantiles and repeated sampling properties of multiple imputation inferences with the model. In Section 5, we present the analysis of the North Carolina lead exposure data. In Section 6, we summarize and suggest research directions. Codes for all analyses are available at https://github.com/jfeldman396/EHQL-Impute.
@@ 2 Gaussian Copula with Nonignorable Missing Data
and to impose some identifying restriction on $p(\boldsymbol{y}^{mis}\mid\boldsymbol{y}^{obs},\boldsymbol{r},\boldsymbol{\gamma})$, also known as the extrapolation model (Linero and Daniels, , 2018). Here, $\boldsymbol{\gamma}$ are parameters of the model for $(\boldsymbol{y},\boldsymbol{r})$. We accomplish these tasks via a Gaussian copula specification, which we now describe.
For each $(\boldsymbol{y}_{i},\boldsymbol{r}_{i})$, let $\boldsymbol{z}_{i}=(\boldsymbol{z}_{\boldsymbol{y}_{i}},\boldsymbol{z}_{\boldsymbol{r}_{i}})$ be a $2p\times 1$ vector of latent variables. Here, $\boldsymbol{z}_{\boldsymbol{y}_{i}}=(z_{i1},\dots,z_{ip})$ and $\boldsymbol{z}_{\boldsymbol{r}_{i}}=(z_{i(p+1)},\dots,z_{i(2p)})$. Let $\boldsymbol{\alpha}=(\mathbf{0}_{p},\alpha_{r_{1}},\dots,\alpha_{r_{p}})$ be a $2p\times 1$ vector, where $\mathbf{0}_{p}$ is a vector of $p$ zeros corresponding to $\boldsymbol{z}_{\boldsymbol{y}_{i}}$ and the next $p$ elements correspond to $\boldsymbol{z}_{\boldsymbol{r}_{i}}$. Let $\boldsymbol{C}_{\boldsymbol{\theta}}$ be a $2p\times 2p$ copula correlation matrix. With generality, we let $\boldsymbol{\theta}$ include parameters that generate a copula correlation; see Section 4 for the specification in our analyses. For $j=1,\dots,p$, let $F_{j}$ be the marginal distribution for $Y_{j}$. To begin, we assume that each $Y_{j}$ is continuous; modifications for discrete $Y_{j}$ are introduced in Section 3. To incorporate unordered categorical variables, we employ the construction in Feldman and Kowal, (2022); see the supplement for details. The data generating model is then
@@
where $\mathbbm{1}_{e}=1$ when the expression $e$ in its index is true and $\mathbbm{1}_{e}=0$ otherwise. The $\boldsymbol{C}_{\boldsymbol{\theta}}$ captures multivariate dependence among the study variables themselves as well as between study variables and nonresponse indicators. The values of $(\alpha_{r_{1}},\dots,\alpha_{r_{p}})$ model the marginal probabilities of missingness. For the study variables, each $z_{ij}$ is transformed to the observed data scale by applying the standard normal cumulative distribution function $\Phi$ and the inverse of the marginal distribution function for $Y_{j}$. For the nonresponse indicators, each $z_{i(p+j)}$ satisfies a probit data augmentation (Chib and Greenberg, , 1998).
For prediction or imputation under copula models with no missing or MCAR data, analysts typically estimate each $F_{j}$ from $\boldsymbol{y}^{obs}$ using the empirical CDF (Hoff, , 2007; Feldman and Kowal, , 2022; Zhao and Udell, , 2020) or with some model (Pitt et al., , 2006; Feldman and Kowal, , 2024). With nonignorable missing data, however, estimates of $\{F_{j}\}_{j=1}^{p}$ computed from $\boldsymbol{y}^{obs}$ can be biased, which in turn affects the quality of imputations of $\boldsymbol{y}^{mis}$. Additionally, inference for $\boldsymbol{C}_{\boldsymbol{\theta}}$ from $(\boldsymbol{y}^{obs},\boldsymbol{r})$ alone may be biased, as we show in Section 4.
We address these potential problems by utilizing some source of auxiliary information about each $F_{j}$. We write this auxiliary information set as $\mathcal{A}_{j}$ and, across study variables, as $\mathcal{A}=\{\mathcal{A}_{j}\}_{j=1}^{p}$. As a first step, we presume $\mathcal{A}_{j}=F_{j}$ for all $j$, i.e., the full marginal distributions are known. In this case, we show that $\boldsymbol{C}_{\boldsymbol{\theta}}$ can be consistently estimated despite the presence of nonignorable missing data (Section 2.1) and that the Gaussian copula implies a version of an AN missingness mechanism (Section 2.2). We leverage these results in Section 3 when $\mathcal{A}$ comprises quantiles rather than full distributions.
In what follows, we include an $R_{j}$ for each $Y_{j}$ in the modeling. When $Y_{j}$ is considered MCAR or has no missing values, analysts can remove $R_{j}$ from (2) and (3), which reduces the dimension of $\boldsymbol{z}_{i}$.
@@ 2.1 Results with Complete Marginals
Let $\boldsymbol{z}_{\boldsymbol{y}}=\{\boldsymbol{z}_{\boldsymbol{y}_{i}}:i=1,\dots,n\}$ and $\boldsymbol{z}_{\boldsymbol{r}}=\{\boldsymbol{z}_{\boldsymbol{r}_{i}}:i=1,\dots,n\}$ be the collections of latent variables for the study variables and nonresponse indicators, respectively. For convenience, we partition $\boldsymbol{z}_{\boldsymbol{y}}=(\boldsymbol{z}^{obs},\boldsymbol{z}^{mis})$ into observed and missing components corresponding to $(\boldsymbol{y}^{obs},\boldsymbol{y}^{mis})$. Define the set restriction $\mathcal{E}(\boldsymbol{r})$ as the condition that $\boldsymbol{z}_{\boldsymbol{r}}$ satisfies the probit constraints in (3) for the realized $\boldsymbol{r}$ across $i=1,\dots,n$ and $j=p+1,\dots,2p$. Because each $F_{j}$ is known from $\mathcal{A}$, the latent variable corresponding to each $y_{ij}^{obs}$ is fixed to $z_{ij}^{obs}=\Phi^{-1}\{F_{j}(y_{ij}^{obs})\}$. Thus, by conditioning on $\{F_{j}\}_{j=1}^{p}$, (1) becomes
@@
With nonignorable missing data, typically one cannot estimate model parameters consistently unless one knows the full distribution of $\boldsymbol{y}^{mis}$. However, as we show in Theorem 1, we need far less information when data follow a copula model: knowledge of the true $\{F_{j}\}_{j=1}^{p}$ is sufficient information to ensure consistent estimation of the copula correlation. For a fixed sample $(\boldsymbol{y}^{obs},\boldsymbol{r})$, the posterior distribution of $(\boldsymbol{C}_{\boldsymbol{\theta}},\boldsymbol{\alpha})$ with known marginals is
@@
The proof is in the supplement. Theorem 1 implies that the copula correlation can be estimated consistently using the observed data $(\boldsymbol{y}^{obs},\boldsymbol{r})$. As a result, we can use the true $\{F_{j}\}_{j=1}^{p}$ and estimates of $\boldsymbol{C}_{\boldsymbol{\theta}}$ to impute $\boldsymbol{y}^{mis}$ from (2)–(3). Furthermore, if $\{F_{j}\}_{j=1}^{p}$ is not known, Theorem 1 suggests that analysts can specify different $\{F_{j}\}_{j=1}^{p}$ to enable interpretable sensitivity analyses. In other words, when the specified $\{F_{j}\}_{j=1}^{p}$ is true, analysts are assured that the model estimates the $\boldsymbol{C}_{0}$ that accords with those marginals consistently.

### v2

@@ Abstract
We present an approach for modeling and imputation of nonignorable missing data. Our approach uses Bayesian data integration to combine (1) a Gaussian copula model for all study variables and missingness indicators, which allows arbitrary marginal distributions, nonignorable missingess, and other dependencies, and (2) auxiliary information in the form of marginal quantiles for some study variables. We prove that, remarkably, one only needs a small set of accurately-specified quantiles to estimate the copula correlation consistently. The remaining marginal distribution functions are inferred nonparametrically and jointly with the copula parameters using an efficient MCMC algorithm. We also characterize the (additive) nonignorable missingness mechanism implied by the copula model. Simulations confirm the effectiveness of this approach for multivariate imputation with nonignorable missing data. We apply the model to analyze associations between lead exposure and end-of-grade test scores for 170,000 North Carolina students. Lead exposure has nonignorable missingness: children with higher exposure are more likely to be measured. We elicit marginal quantiles for lead exposure using statistics provided by the Centers for Disease Control and Prevention. Multiple imputation inferences under our model support stronger, more adverse associations between lead exposure and educational outcomes relative to complete case and missing-at-random analyses.
Keywords: Bayesian inference; imputation; MNAR; data integration; nonresponse
@@ 1 Introduction
The Gaussian copula is a flexible joint distribution for multivariate data. The model can characterize complex dependencies while also capturing non-Gaussian marginal distributions. Methodological advances have made the Gaussian copula compatible with mixed data types (Hoff, 2007; Feldman and Kowal, 2022), increased its scalability to high-dimensional variable sets (Murray et al., 2013), and improved its ability to capture non-linearity and interactions (Feldman and Kowal, 2024). Because of these appealing features, it has been deployed in numerous applications, for example, in economics and finance (Fan and Patton, 2014), marketing and management (Eckert and Hohberger, 2023), and political science (Chiba et al., 2015).
@@
The Gaussian copula can be readily implemented when data have missing values. Indeed, researchers (e.g., Käärik and Käärik, 2010; Di Lascio et al., 2015; Zhao and Udell, 2020; Hollenbach et al., 2021; Hoff, 2022; Christoffersen et al., 2023) have suggested Gaussian copulas as multivariate, joint modeling engines for (multiple) imputation. Existing methods, however, tend to assume that missingness among the study variables is missing completely at random (MCAR) or missing at random (MAR) (Rubin, 1976). We are not aware of methodology for itemwise nonignorable missingness with Gaussian copula models.
@@
In this article, we develop a framework to handle nonignorable missing data in Gaussian copula models. Our general strategy is to specify a Bayesian Gaussian copula model that includes both the study variables and the missingness indicators, coupled with auxiliary information on the marginal distributions of the study variables that we integrate directly into the model. Using this strategy, we prove that analysts need to know only a small, arbitrary set of auxiliary marginal quantiles to enable consistent estimation of the copula correlation; entire distributions are not necessary. Further, we show that the copula model encodes a specific identifying assumption about the missingness mechanism, namely that it is a type of additive nonignorable (AN) mechanism (Hirano et al., 2001; Sadinle and Reiter, 2019). We develop algorithms for estimating the copula correlation that result in significant computational gains relative to alternative copula models, which allows fitting to scale to data with (moderately) large sample sizes in reasonable time using typical computational setups. We also present strategies for estimating other quantiles of the marginal distributions beyond those in the auxiliary information. Our approach provides uncertainty quantification for the unknown marginals, which we utilize for multiple imputation (Rubin, 1987).
@@
We develop the methodology for settings where marginal quantiles for variables subject to missingness are available in external data sources. For example, previewing the application in Section 5, suppose an analysis involving health measurements suffers from nonignorable missingness, e.g., people likely to have unconcerning values are not measured. For many health measurements, there exists information on marginal quantiles from national benchmark surveys or administrative databases. We seek to use the auxiliary quantiles to adjust for nonignorable missingness in the analysis at hand. When auxiliary quantiles are not available or known precisely, analysts can posit different plausible marginal quantiles and assess the sensitivity of ultimate analyses under those specifications, e.g., via multiple imputation analyses. We note that the use of known marginal distributions for imputation of nonignorable missing data also has been proposed for categorical data models (Pham et al., 2018; Akande et al., 2021; Deng et al., 2013; Si et al., 2015, 2016; Tang et al., 2024), but not for continuous and mixed variables like we do here.
@@
We apply the methodology in data comprising health, socioeconomic, demographic and educational measurements collected on over 170,000 children residing in the state of North Carolina. Among the study variables are end-of-grade math and reading test scores and a measure of lead exposure, the latter of which is subject to abundant missingness that is likely nonignorable. The state requires children at high risk of lead exposure to be measured, but not children at low risk of lead exposure. To inform the imputation of the missing values, we leverage marginal quantiles on lead exposure published by the U.S. Centers for Disease Control and Prevention. Using this information, we find lead exposure apparently is more adversely associated with test scores than suggested from complete cases or MAR analyses.
@@
The remainder of this article is organized as follows. In Section 2, we present the Gaussian copula model for nonignorable missing data which leverages information on marginal distributions. Here we present two theoretical results, namely (i) that fully specified marginal distributions provide posterior consistency of the copula correlation in the presence of nonignorable missingness, and (ii) a derivation that the Gaussian copula implies a version of the AN missingness mechanism. In Section 3, we modify the copula model and develop imputation strategies for settings where the auxiliary marginal information constitutes an incomplete set of marginal quantiles. Here we present our main theorem: under the AN missingness mechanism, this limited auxiliary information still allows consistent estimation of the copula correlation. In Section 4, we present simulations studying the effect of differing amounts of auxiliary quantiles and repeated sampling properties of multiple imputation inferences with the model. In Section 5, we present the analysis of the North Carolina lead exposure data. In Section 6, we summarize and suggest research directions. Supplementary material includes proofs of all theorems and additional results from simulations and the North Carolina data analysis. Codes for all analyses are available at https://github.com/jfeldman396/EHQL-Impute.
@@ 2 Gaussian Copula with Nonignorable Missing Data
and to impose some identifying restriction on $p(\boldsymbol{y}^{mis}\mid\boldsymbol{y}^{obs},\boldsymbol{r},\boldsymbol{\gamma})$, also known as the extrapolation model (Linero and Daniels, 2018). Here, $\boldsymbol{\gamma}$ are parameters of the model for $(\boldsymbol{y},\boldsymbol{r})$. We accomplish these tasks via a Gaussian copula specification, which we now describe.
For each $(\boldsymbol{y}_{i},\boldsymbol{r}_{i})$, let $\boldsymbol{z}_{i}=(\boldsymbol{z}_{\boldsymbol{y}_{i}},\boldsymbol{z}_{\boldsymbol{r}_{i}})$ be a $2p\times 1$ vector of latent variables. Here, $\boldsymbol{z}_{\boldsymbol{y}_{i}}=(z_{i1},\dots,z_{ip})$ and $\boldsymbol{z}_{\boldsymbol{r}_{i}}=(z_{i(p+1)},\dots,z_{i(2p)})$. Let $\boldsymbol{\alpha}=(\mathbf{0}_{p},\alpha_{r_{1}},\dots,\alpha_{r_{p}})$ be a $2p\times 1$ vector, where $\mathbf{0}_{p}$ is a vector of $p$ zeros corresponding to $\boldsymbol{z}_{\boldsymbol{y}_{i}}$ and the next $p$ elements correspond to $\boldsymbol{z}_{\boldsymbol{r}_{i}}$. Let $\boldsymbol{C}_{\boldsymbol{\theta}}$ be a $2p\times 2p$ copula correlation matrix. With generality, we let $\boldsymbol{\theta}$ include parameters that generate a copula correlation; see Section 4 for the specification in our analyses. For $j=1,\dots,p$, let $F_{j}$ be the marginal distribution for $Y_{j}$. To begin, we assume that each $Y_{j}$ is continuous. Modifications for discrete $Y_{j}$ are introduced in Section 3. The data generating model is then
@@
where $\mathbbm{1}_{e}=1$ when the expression $e$ in its index is true and $\mathbbm{1}_{e}=0$ otherwise. The $\boldsymbol{C}_{\boldsymbol{\theta}}$ captures multivariate dependence among the study variables themselves as well as between study variables and nonresponse indicators. The latter correlations can encode potentially nonignorable missingness, as discussed in Section 2.2. The values of $(\alpha_{r_{1}},\dots,\alpha_{r_{p}})$ model the marginal probabilities of missingness. For the study variables, each $z_{ij}$ is transformed to the observed data scale by applying the standard normal cumulative distribution function $\Phi$ and the inverse of the marginal distribution function for $Y_{j}$. For the nonresponse indicators, each $z_{i(p+j)}$ satisfies a probit data augmentation (Chib and Greenberg, 1998).
For prediction or imputation under copula models with no missing or MCAR data, analysts typically estimate each $F_{j}$ from $\boldsymbol{y}^{obs}$ using the empirical CDF (Hoff, 2007; Feldman and Kowal, 2022; Zhao and Udell, 2020) or with some model (Pitt et al., 2006; Feldman and Kowal, 2024). With nonignorable missing data, however, estimates of $\{F_{j}\}_{j=1}^{p}$ computed from $\boldsymbol{y}^{obs}$ can be biased, which in turn affects the quality of imputations of $\boldsymbol{y}^{mis}$. Additionally, inference for $\boldsymbol{C}_{\boldsymbol{\theta}}$ from $(\boldsymbol{y}^{obs},\boldsymbol{r})$ alone may be biased (Section 4.1).
We address these potential problems by utilizing auxiliary information about each $F_{j}$. We write this auxiliary information set as $\mathcal{A}_{j}$ and, across study variables, as $\mathcal{A}=\{\mathcal{A}_{j}\}_{j=1}^{p}$. As a first step, we presume $\mathcal{A}_{j}=F_{j}$ for all $j$, i.e., the full marginal distributions are known. In this case, we show that $\boldsymbol{C}_{\boldsymbol{\theta}}$ can be consistently estimated despite the presence of nonignorable missing data (Section 2.1) and that the Gaussian copula implies a version of an AN missingness mechanism (Section 2.2). We leverage these results in Section 3 when $\mathcal{A}$ comprises incomplete information on the marginal distributions.
In what follows, we include a $R_{j}$ for each $Y_{j}$ in the modeling. When $Y_{j}$ is considered MCAR or has no missing values, analysts can remove $R_{j}$ from (2) and (3), which reduces the dimension of $\boldsymbol{z}_{i}$.
@@ 2.1 Results with Complete Marginals
Let $\boldsymbol{z}_{\boldsymbol{y}}=\{\boldsymbol{z}_{\boldsymbol{y}_{i}}:i=1,\dots,n\}$ and $\boldsymbol{z}_{\boldsymbol{r}}=\{\boldsymbol{z}_{\boldsymbol{r}_{i}}:i=1,\dots,n\}$ be the latent variables for the study variables and nonresponse indicators, respectively. For convenience, we partition $\boldsymbol{z}_{\boldsymbol{y}}=(\boldsymbol{z}^{obs},\boldsymbol{z}^{mis})$ into observed and missing components corresponding to $(\boldsymbol{y}^{obs},\boldsymbol{y}^{mis})$. Define the set restriction $\mathcal{E}(\boldsymbol{r})$ as the condition that $\boldsymbol{z}_{\boldsymbol{r}}$ satisfies the probit constraints in (3) for the realized $\boldsymbol{r}$ across $i=1,\dots,n$ and $j=p+1,\dots,2p$. Because each $F_{j}$ is known from $\mathcal{A}$, the latent variable corresponding to each $y_{ij}^{obs}$ is fixed to $z_{ij}^{obs}=\Phi^{-1}\{F_{j}(y_{ij}^{obs})\}$. Thus, by conditioning on $\{F_{j}\}_{j=1}^{p}$, (1) becomes
@@
With nonignorable missing data, typically one cannot estimate model parameters consistently unless one knows the full distribution of $\boldsymbol{y}^{mis}$. However, as stated formally in Theorem 1, when the study variables and missingness indicators follow a copula model, knowledge of the true $\{F_{j}\}_{j=1}^{p}$ is sufficient information to ensure consistent estimation of the copula correlation. For a fixed sample $(\boldsymbol{y}^{obs},\boldsymbol{r})$, the posterior distribution of $(\boldsymbol{C}_{\boldsymbol{\theta}},\boldsymbol{\alpha})$ with known marginals is
@@
Theorem 1 implies that $\boldsymbol{C}_{0}$ can be estimated consistently using the observed data $(\boldsymbol{y}^{obs},\boldsymbol{r})$ and the true $\{F_{j}\}_{j=1}^{p}$. With these quantities, we are able to impute $\boldsymbol{y}^{mis}$ from (2)–(3). In practice, the full $\{F_{j}\}_{j=1}^{p}$ may not be known. In this case, Theorem 1 suggests that analysts can specify different $\{F_{j}\}_{j=1}^{p}$ to enable interpretable sensitivity analyses. However, as we show in Section 3, analysts need far less information about the marginals for consistent estimation of $\boldsymbol{C}_{0}$ and subsequent implementation of multiple imputation.